Let T be the linear operator on R


3 defined by


T(x1, x2, x3) = (3x1, x1 − x2 , 2x1 + x2 + x3 )

Is T invertible? If so, find a rule for T −1 like the one which defines T.


1
Expert's answer
2021-01-04T20:18:56-0500

Given that ,

"T:\R^3 \rightarrow \R^3" is a linear transformation defined by

"T(x_1,x_2,x_3)=(3x_1,x_1-x_2,2x_1+x_2+x_3)"


"\therefore" We can define , "T=\begin{pmatrix}
3 &0&0 \\
1&-1&0\\
2&1&1
\end{pmatrix}" with respect to usual basis .


"\therefore Det(T)=\begin{vmatrix}
3&0&0 \\
1&-1&0\\
2&1&1
\end{vmatrix}" "=-3"


"\therefore T" is invertible .

Now , we have to find "T^{-1}" .

For "T^{-1}" , first we have to find adj("T" ).


"adj(T)=\begin{pmatrix}
-1&-1 &3 \\
0&3&-3 \\
0&0&-3
\end{pmatrix}" .


"\therefore T^{-1}=\frac{(adjT)'}{Det(T)}" "=\frac{1}{-3} \begin{pmatrix}
-1&0&0\\
-1&3&0 \\
3&-3&-3
\end{pmatrix}"


"=\begin{pmatrix}
\frac{1}{3}&0&0 \\
\frac{1}{3}&-1&0 \\
-1&1&1
\end{pmatrix}"


"\therefore T^{-1}:\R^3\rightarrow \R^3" defined by

"T^{-1}(x_1,x_2,x_3)=(\frac{x_1}{3},\frac{x_1}{3}-x_2,-x_1+x_2+x_3)"